OpenStudy (anonymous):

how to you express cos4x in terms of sin

OpenStudy (anonymous):

First find cos(2x). cos(2x) = cos(x+x) = cosxcosx - sinxsinx = cos^2x - sin^2x =1-2sin^2(x) Then cos(4x) =cos(2x+2x) = cos2xcos2x - sin2xsin2x = (1-2sin^2x)(1-2sin^2x) - 2sinxcosx.2sinxcosx = (1-4sin^2x +4sin^4x) -4sin^2xcos^2x = 1-4sin^2x+4sin^4x -4sin^2x(1-sin^2x) = 1-8sin^2(x)+8sin^4(x)

OpenStudy (anonymous):

Thanks